# The Case For Black Hole Remnants: A Review

### v1

@@ I Introduction: The Puzzles of Black Hole Evaporation
Hawking showed that black holes can radiate quantum mechanically. For an asymptotically flat Schwarzschild black hole, the Hawking temperature [^1] We will use the units $G=\hbar=c=4\pi\epsilon_{0}=k_{B}=1$. is given by $T=1/8\pi M$. This is inversely proportional to the black hole mass $M$. Therefore, as the black hole loses mass and becomes smaller, it also becomes hotter. The temperature is not bounded above and so diverges in the limit $M\to 0$. The obvious question is: what is the end state of Hawking evaporation? The usual model of Hawking evaporation is governed by the simple ODE of the form
@@
where $c$ is some constant related to the particle species, $A$ the area of the black hole and $T$ the temperature. The right hand side is just Stefan-Boltzmann law for blackbody emission. The cross sectional area is not simply $A$ as it should be determined by the gray-body factor, and even in the geometric optical limit the length scale is determined by the impact parameter of the photon sphere instead of just the Schwarzschild radius. Still, modulo these nitty-gritty details, the qualitative feature can be understood by Eq.(1). Since $T\sim 1/M$ and $A\sim M^{2}$, we see that $dM/dt\sim-M^{-2}$, which we can solve to find that a black hole whose initial mass is $M$ will evaporate completely in a finite time that scales as $\sim M^{3}$. In general relativity, black holes have “no hair”, which means that there is no other physical parameters that characterize a black hole other than the mass, electric (and hypothetically, magnetic) charge, and angular momentum. Including charge and angular momentum changes the history of the evolution somewhat, but since angular momentum and charge are always radiated away more efficiently than mass loss (in the standard picture of Hawking evaporation), the end state is always a Schwarzschild black hole, so it suffices to consider that as far as remnants are concerned.
@@
A model is of course, only as good as the assumptions we put into it. The above conclusion of a complete evaporation is based on the validity of Eq.(1), i.e., on the validity of the Stefan-Boltzmann law. Indeed, it is commonly accepted that at sufficiently high energy scale (i.e., high temperature, i.e., small black holes), new physics would likely kick in. This could happen before the Planck scale, and around the Planck scale we would expect quantum gravity effects to become important. Therefore, this already means that the end state of black hole evaporation is not well understood, and on this ground there is no reason to prefer a complete evaporation over the possibility that evaporation somehow effectively stops, becoming a “remnant”. By remnant, we mean a state of black holes at late time of its evolution, whose lifetime can be infinite, or finite but longer than $M^{3}$, say $M^{4}$.
@@
The aim of this review is to provide an overview of black hole remnants that is accessible to colleagues who are familiar with black holes but know little about remnants. There is already a more comprehensive review which I co-authored with Yeom and Chen almost 10 years ago [1], and readers are encouraged to refer to that for more details. In this review, I will instead mentioned some new progress and insights that have been made since then, focusing on (1) how remnants can possibly form and (2) traditional criticisms against remnants and why they do not rule out remnants. Finally we will end with some discussions.
@@
The possibility of a black hole remnant was proposed already at least as far back as 1987, in the work of Aharonov, Casher, and Nussinov [2], on the ground that Hawking evaporation should satisfy unitarity of quantum evolution [^2] Essentially, the problem is that if a pure state forms a black hole, while the entirety of quantum state remains pure, the interior and exterior configurations are mixed. Since the interior states presumably cannot escape the black hole, they somehow “disappear” if the black hole completely evaporates. This implies that we have a pure-to-mixed evolution, violating unitarity.. They called remnants “Planckon”. This would be deemed “impossible” by later works that instead prefer to maintain unitarity by extracting quantum information from the Hawking radiation itself. In order to better appreciate the issue, we now give a quick summary of the infamous “information paradox” and the “firewall” or AMPS paradox [3, 4] that came after it.
@@
However, the AMPS paradox pointed out that doing so would lead to another curious issue, namely that at late time (after the black hole entropy has been halved, called the Page time [5, 6, 7]), the black hole horizon ceases to be “uneventful” and would instead be replaced by a high energy region, dubbed a “firewall”. Any observer that falls through a black hole in classical general relativity would just pass right through the horizon (a mathematical surface!), but once the firewall sets in, an in-falling observer would now be incinerated at the firewall. The reason for the existence of the firewall is as follows. Hawking particles are formed by pair-production from the vacuum “near” [^3] This is not completely correct. For a Schwarzschild black hole the particles can be emitted in a diffused region called “quantum atmosphere” that extends some $O(1)$ to $O(10)$ distance away from the horizon [8]. the horizon, one of which falls into the black hole while the other is emitted to infinity. The usual argument is that (1) the pair-produced particles are maximally entangled, and that (2) information only starts to be released after Page time. The Hawking particles emitted after Page time needs to purify the earlier Hawking radiation, and therefore they cannot be maximally entangled with their own partner that falls into the black hole. This is due to the monogamy of quantum information: a particle cannot be maximally entangled with two particles at the same time. Without the maximal entanglement between the outgoing and ingoing pairs, the field configuration across the horizon becomes discontinuous. This in turn means that the Hamiltonian (energy) of the field diverges across the horizon (since it contains terms like $\partial^{2}\phi$, that involves the derivative of the field) – the firewall. Such a drastic departure from general relativity [^4] The situation is even worse than initially expected. If firewalls can in fact exist, then there is a nonzero probability that a firewall can occur arbitrarily far away from the horizon, a “naked firewall” [9]. Due to the divergence in energy density, a firewall is essentially a singularity; a naked firewall is therefore as bad as a naked singularity. We will have more discussions later on naked singularity. led to a fierce debate in the community as to what has possibly gone wrong in the analysis.
@@
Indeed, there are assumptions that go into the AMPS paradox, and “resolutions” of the paradox by challenging these premises can be found throughout the literature. For example, maximal entanglement has been challenged [10, 11]. For our purpose, however, it is the “central dogma” [12] – the claim that a black hole behaves like a quantum system whose number of degrees of freedom is proportional to the horizon area – that requires a closer examination (see, e.g., [13] for some criticisms of the dogma). The dogma is related to known results that there exist upper bound for how much information one can squeeze into a finite size region with a given energy, namely the Bekenstein bound [14] ($S\leqslant 2\pi RE$) and its various more recent generalizations. Thus, if the central dogma holds, and if Hawking evaporation does not contain any information, then problems already arise when a black hole becomes sufficiently small (after Page time), regardless of whether there is a complete evaporation at the end. This is because the horizon would eventually be too small to bound the large amount of information inside. In particular this would mean that the remnant picture is incorrect. However, as we will discuss later on, it could be that the central dogma is not correct, and the remnant picture may still well be viable. But for now, let us begin by explaining how one can obtain remnants in various quantum gravitational models.
@@ II How to Obtain Remnants
A black hole becomes a remnant when it stops evaporating. For example, this can happen if it becomes an extremal black hole. In this context, by an extremal black hole we mean the temperature of the black hole is zero (these two notions do not always coincide). Note that there is no need for it to become exactly extremal: just by asymptotically approaching an extremal state the black hole evaporation time scale becomes indefinitely long. Zero temperature extremal black holes are black holes whose inner horizon coincides with the outer horizon. In classical general relativity, one needs to include either a charge or angular momentum in order for a black hole to have two horizons. However, if one envisages removing the black hole singularity, then it is common to end up with “regular black holes” with two horizons, even without charge or angular momentum. Note, however, inner horizons are themselves highly unstable and would likely collapse into another singularity. This seems to defeat the whole point of regularizing the singularity in the first place. Worse, the end state of this instability is unknown, so we cannot say whether there will still be a remnant afterward. One recent progress is to construct black holes with “inner extremal horizon”, with zero surface gravity, and therefore avoids the classical blueshift instability. Unfortunately this does not seem to avoid the Aretakis instability (classical perturbation does not decay) [15, 16, 17, 18]. In addition, semi-classically the inner extremal horizon remains unstable [19]. This way of obtaining a remnant is therefore quite challenging conceptually, but clearly more studies are required, as the properties of inner horizon of a black hole are still unclear in general [20], and the possibility that black holes can asymptote towards some extremal configurations and thus become an effective remnant cannot yet be ruled out [21]. Note that extremal black hole is not the only possible end state, as some regular black holes can evolve into a horizonless object [22, 23, 24].
@@
Another possibility is to consider quantum gravitational effect that kicks in when the black hole is sufficiently small. Note the subtle difference: in the contexts of regular black holes, while the classical singularity could be removed by some kind of quantum gravity effect, remnants are the result of the existence of extremal state. The inner horizon — which would not be there without the quantum gravitational effect regularizing the horizon — exists already when the black hole is still astronomically large. Thus, these extremal remnants are not directly caused by quantum gravitational correction, only indirectly so. A more direct influence from quantum gravity can happen at or near the Planck scale, without necessarily modifying the black hole solution when $M$ is still large, and this can be studied phenomenologically by employing the “generalized uncertainty principle” (GUP).
@@
where $\alpha$ is a dimensionless parameter, typically taken to be of order unity. We will first discuss the case in which $\alpha>0$. From Eq.(2) we can obtain the following inequality:
@@
Hawking temperature is modified under GUP. For an asymptotically flat Schwarzschild black hole this modification was obtained heuristically in [25] by considering the characteristic energy of the Hawking particle as $E=pc$ and by identifying $\Delta x$ with the horizon scale $\Delta x=2GM/c^{2}$. This gives (upon multiplying with a normalization constant prefactor $1/2\pi$ in order to obtain the correct temperature when $\alpha\to 0$):
@@
The problem with such a heuristic treatment is that we cannot be sure of its validity. After all, while the method works for the Schwarzschild case, it does not lead to the correct known temperature for the Kerr or Reissner-Nordström black hole [26], so how could we be confident that it works when there is a GUP correction? For other criticisms for heuristic treatments of GUP, see [27, 28, 29].
@@
where $M_{p}$ denotes the Planck mass. This could be a sign that modifying the Hawking temperature without modifying the Schwarzschild metric may be problematic. Of course, the whole idea of a classical, well-defined, metric tensor might already fail near the Planck scale, so it is not clear if any modified metric is trustworthy. Nevertheless, we note that the entropy of the black hole, which can be obtained by integrating the first law of black hole thermodynamics, can be shown to satisfy (returning to Planck units hereinafter)
@@
Surprisingly, the problem of the final non-evaporating but nonzero temperature black hole can be avoided. This is due to another effect of GUP that can prevent black hole evaporation even when the black hole is still astronomically large (and therefore when the temperature remains largely unmodified), just after a scrambling time [30]. If correct, this renders the aforementioned Planckian size black hole remnant irrelevant except for black holes that were formed to be microscopic in the beginning. For a Schwarzschild black hole, the scrambling time [31, 32] is $r_{S}\log(r_{s}/L_{p})$, where $r_{S}=2M$ denotes the horizon; it is the time scale for information to be scrambled over the horizon after being dropped into a black hole. Note that this time scale is far shorter than $\sim r_{S}^{3}$, the lifetime of a black hole. Thus, if a large black hole stops evaporating after the scrambling time, then it becomes a macroscopic remnant and the peculiarity mentioned above never arises. Intuitively this result can be appreciated as follows. First, we recall the heuristic derivation of Hawking temperature above, in which the Hawking temperature scale is roughly set by $\Delta p$, and by the uncertainty principle $\Delta x\sim 1/\Delta p$, we can trace $\Delta x$ back to the horizon scale, i.e., Hawking radiation comes from the region around the black hole. However, with GUP, it can be shown that beyond the scrambling time, $\Delta p\gg\alpha/L_{p}^{2}$, which in turn implies that $\Delta x\gg r_{S}$, which cannot be interpreted as particle creation by the black hole [30]. See also [33], in which the same conclusion was reached using a string theoretic approach. Note that if correct, this also avoids the firewall problem since the black hole never becomes small enough to get pass the Page time scale. The remnant is essentially a classical black hole. To contrast with this scrambling time scale remnant, in another recent approach, it is argued that black hole stops evaporating after it has emitted about half of its initial mass, due to “memory burden effect” [34], in which a system with high capacity of information storage (such as a black hole) is stabilized by the load of information it carries. Thus we see that it is not necessarily the case that remnants are microscopic.
@@
It is almost poetic that a quantum gravitational correction like the one in GUP, which we usually think can only be important when the black hole is small, can give rise to a drastic deviation from semi-classical picture even when the black hole is still huge (and even renders the black hole effectively classical with no Hawking radiation). Yet, this is not the first time we have seen such an unexpected behavior. For another example, consider the Chandrasekhar limit of a white dwarf star. White dwarfs are stable because electron degenerate pressure counteracts gravitational collapse. This pressure is inherently quantum mechanical in nature, it arises from the jittering motion (large $\Delta p$) of the particles inside the star, which increases as the star becomes smaller (small $\Delta x$). Once GUP is taken into account, it seems that the Chandrasekhar limit is destroyed [35, 36]. Dynamical stability analysis can restore an upper mass limit for white dwarfs for a certain range of the GUP parameter [37], which explains why in astrophysical observation we do not observe arbitrarily large white dwarfs. Still, here too we see that a supposedly small quantum gravity effect can lead to big changes to physics well outside the quantum gravity regime.
@@
In [38] I considered Hawking evaporation under the assumption that the GUP parameter could be negative. In such a scenario there is no longer a minimum mass, and so the black hole can continue to evaporate indefinitely. However, its lifetime turns out to be infinite, so we still have a remnant that asymptotes to zero mass. This avoids another peculiar property: the zero mass black hole (or rather, no more black hole!) has a finite nonzero temperature. See also [39] for more discussions.
@@
So far we have discussed extremal black holes and the GUP approaches that led to remnants. In recent years, another promising possibility has come to light: a white hole remnant, which is a scenario supported by loop quantum gravity [40, 41, 42, 43, 44, 45, 46], see the recent review [47] for details. The basic idea is simple: the static Schwarzschild solution does not distinguish the direction of time, and indeed in the maximally extended spacetime we see that there are black hole and white hole, described by the same metric tensor. Classically, we do not usually care about the white hole part because it is unphysical. However, quantum mechanically one should consider a superposition of both black and white holes, and the possible tunneling from the black hole state to the white hole one. This probability becomes significant as the black hole becomes sufficiently small. In this scenario, a black hole with initial mass $M$ evaporates with lifetime that scales like $M^{3}$ as in the standard Hawking picture. However, before it completely evaporates away, it tunnels into the white hole state, which then become a remnant that slowly radiates away. The time scale for this diffuse emission is about $M^{4}$ [46]. Such a white hole remnant does not live indefinitely. However, like the more traditional black hole remnant it can conceivably still solve the information paradox, as we will discuss in the next section.

### v2

@@ I Introduction: The Puzzles of Black Hole Evaporation
Classical black holes are “black” in the sense that nothing, not even light, can escape from inside the horizon. Anything that falls into a black hole is forever lost to the outside world. This naive picture is drastically changed after Hawking showed that black holes can radiate quantum mechanically. For an asymptotically flat Schwarzschild black hole, the Hawking temperature [^1] We will use the units $G=\hbar=c=4\pi\epsilon_{0}=k_{B}=1$. is given by $T=1/8\pi M$. This is inversely proportional to the black hole mass $M$. Therefore, as the black hole loses mass and becomes smaller, it also becomes hotter. The temperature is not bounded above and so diverges in the limit $M\to 0$. The obvious question is: what is the end state of Hawking evaporation? The usual model of Hawking evaporation is governed by the simple ODE of the form
@@
where $C$ is some constant related to the particle species, $A$ the area of the black hole and $T$ the temperature. The right hand side is just the Stefan-Boltzmann law for blackbody emission. The cross sectional area is not simply $A$ as it should be determined by the gray-body factor, and even in the geometric optical limit the length scale is determined by the impact parameter of the photon sphere instead of just the Schwarzschild radius. Still, modulo these nitty-gritty details, the qualitative feature can be understood by Eq.(1). Since $T\sim 1/M$ and $A\sim M^{2}$, we see that $dM/dt\sim-M^{-2}$, which we can solve to find that a black hole whose initial mass is $M$ will evaporate completely in a finite time that scales as $t_{\text{life}}\sim M^{3}$. In general relativity, black holes have “no hair”, which means that there is no other physical parameters that characterize a black hole other than the mass, electric (and hypothetically, magnetic) charge, and angular momentum. Including charge and angular momentum changes the history of the evolution somewhat, but since angular momentum and charge are always radiated away more efficiently than mass loss (in the standard picture of Hawking evaporation), the end state is always a Schwarzschild black hole, so it suffices to consider that as far as remnants are concerned.
@@
A model is of course, only as good as the assumptions we put into it. The above conclusion of a complete evaporation is based on the validity of Eq.(1), i.e., on the validity of the Stefan-Boltzmann law. Indeed, it is commonly accepted that at sufficiently high energy scale (i.e., high temperature; i.e., small black holes), new physics would likely kick in. This could happen before the Planck scale, and around the Planck scale we would expect quantum gravity effects to become important. Therefore, this already means that the end state of black hole evaporation is not well understood, and on this ground there is no reason to prefer a complete evaporation over the possibility that evaporation somehow effectively stops, becoming a “remnant”. By remnant, we mean a state of black holes at late time of its evolution, whose lifetime can be infinite, or finite but longer than $M^{3}$, say $M^{4}$.
@@
The aim of this review is to provide an overview of black hole remnants that is accessible to colleagues and students who are familiar with black holes but know little about remnants. There is already a more comprehensive review which I co-authored with Yeom and Chen almost 10 years ago [1], and readers are encouraged to refer to that for more details. In this review, I will instead mention some new progress and insights that have been made since then, focusing on (1) how remnants can possibly form and (2) traditional criticisms against remnants and why they do not rule out remnants. Finally we will end with some discussions.
@@
The possibility of a black hole remnant was proposed already at least as far back as 1987, in the work of Aharonov, Casher, and Nussinov [2], on the ground that Hawking evaporation should satisfy unitarity of quantum evolution [^2] Essentially, the problem is that if a pure state forms a black hole, while the entirety of quantum state remains pure, the interior and exterior configurations are mixed. Since the interior states presumably cannot escape the black hole, they somehow “disappear” if the black hole completely evaporates. This implies that we have a pure-to-mixed evolution, violating unitarity.. They called remnants “Planckons”. Indeed, microscopic black holes may look like an “elementary particle” at least from the view of an exterior observer (the interior of a small black hole may still be macroscopically large; see below). Some authors have interpreted a black hole to be a collection of many microscopic “black hole atoms” [3]. However, the remnant scenario would be deemed “impossible” by later works that instead prefer to maintain unitarity by extracting quantum information from the Hawking radiation itself (see, e.g., the review [4]). In order to better appreciate the issue, we now give a quick summary of the infamous “information paradox” and the “firewall” or AMPS paradox [5, 6] that came after it.
@@
However, the AMPS paradox pointed out that doing so would lead to another curious issue, namely that at late time (after the black hole entropy has been halved, called the Page time [7, 8, 9]), the black hole horizon ceases to be “uneventful” and would instead be replaced by a high energy region, dubbed a “firewall”. Any observer that falls through a black hole in classical general relativity would just pass right through the horizon (a mathematical surface!), but once the firewall sets in, an in-falling observer would now be incinerated at the firewall. The reason for the existence of the firewall is as follows. Hawking particles are formed by pair-production from the vacuum “near” [^3] This is not completely correct. For a Schwarzschild black hole the particles can be emitted in a diffused region called “quantum atmosphere” that extends some $O(1)$ to $O(10)$ distance away from the horizon [10]. the horizon, one of which falls into the black hole while the other is emitted to infinity. The usual argument is that (1) the pair-produced particles are maximally entangled, and that (2) information only starts to be released after Page time. The Hawking particles emitted after Page time needs to purify the earlier Hawking radiation, and therefore they cannot be maximally entangled with their own partner that falls into the black hole. This is due to the monogamy of quantum information: a particle cannot be maximally entangled with two particles at the same time. Without the maximal entanglement between the outgoing and ingoing pairs, the field configuration across the horizon becomes discontinuous. This in turn means that the Hamiltonian (energy) of the field diverges across the horizon (since it contains terms like $\partial^{2}\phi$, that involves the derivative of the field); this is the firewall. An observer that falls into a sufficiently old black hole would encounter the firewall and be destroyed (astrophysical black holes are not yet old enough). Such a drastic departure from the “no drama” scenario of general relativity [^4] The situation is even worse than initially expected. If firewalls can in fact exist, then there is a nonzero probability that a firewall can occur arbitrarily far away from the horizon, a “naked firewall” [11]. Due to the divergence in energy density, a firewall is essentially a singularity; a naked firewall is therefore as bad as a naked singularity. We will have more discussions later on naked singularity. led to a fierce debate in the community as to what has possibly gone wrong in the analysis.
@@
Indeed, there are assumptions that go into the AMPS paradox, and “resolutions” of the paradox by challenging these premises can be found throughout the literature. For example, the requirement for maximal entanglement has been challenged [12, 13]. For our purpose, however, it is the “central dogma” [14] — the claim that a black hole behaves like a quantum system whose number of degrees of freedom is proportional to the horizon area — that requires a closer examination (see, e.g., [15] for some criticisms of the dogma). The dogma is related to known results that there exists an upper bound for how much information one can squeeze into a finite size region with a given energy, namely the Bekenstein bound [16] ($S\leqslant 2\pi RE$) and its various more recent generalizations. Thus, if the central dogma holds, and if Hawking evaporation does not contain any information, then problems already arise when a black hole becomes sufficiently small (after Page time), regardless of whether there is a complete evaporation at the end. This is because the horizon would eventually be too small to bound the large amount of information inside. In particular this would mean that the remnant picture for solving the information paradox is incorrect [^5] Note that remnants themselves are not incompatible with the central dogma in the sense that if the latter is correct, then a microscopic remnant can only have a tiny amount of information, and therefore cannot solve the information paradox. For the same reason, the firewall would arise as in the standard picture; see [1]. On the other hand, this also means that if one’s goal is to solve the information paradox, it is not enough to just come up with a model in which there is a remnant at the end of the evaporation — one has to assume the failure of the central dogma independently of the existence of the remnant. It is this assumption that is at the heart of the debate.. However, as we will discuss later on, it could be that the central dogma is not correct, and the remnant picture may still well be viable. But for now, let us begin by explaining how one can obtain remnants in various quantum gravitational models.
@@ II How to Obtain Remnants
A black hole becomes a remnant when it stops evaporating. For example, this can happen if it becomes an extremal black hole. In this context, by an extremal black hole we mean the temperature of the black hole is zero (these two notions do not always coincide). Note that there is no need for it to become exactly extremal: just by asymptotically approaching an extremal state the black hole evaporation time scale becomes indefinitely long. Zero temperature extremal black holes are black holes whose inner horizon coincides with the outer horizon. In classical general relativity, one needs to include either a charge or angular momentum in order for a black hole to have two horizons. However, if one envisages removing the black hole singularity, then it is common to end up with “regular black holes” with two horizons [^6] This is not necessarily the case. For example, in [17], the authors constructed a regular black hole by considering a massive source in a quantum superposition of locations. The black hole only has a single horizon, with the zero-temperature extremal case separating the black hole configuration with a (horizonless) wormhole configuration. For another example, see [18], in which a single-horizon regular black hole whose singularity is replaced by a bounce was constructed via an effective polymerization scheme in the Loop Quantum Gravity approach., even without charge or angular momentum. Such black holes can then radiate indefinitely as they approach the zero temperature limit (whereas for the charged and/or rotating case in general relativity, the charge and angular momentum are typically shed faster than the mass and so the final state is still the Schwarzschild solution). Bardeen black hole [19], non-commutative geometry black hole [20], asymptotically safe gravity black hole [21], and two-dimensional dilaton black hole [22] are just some examples.
Note, however, inner horizons are themselves highly unstable and would likely collapse into another singularity. This seems to defeat the whole point of regularizing the singularity in the first place. Worse, the end state of this instability is unknown, so we cannot say whether there will still be a remnant afterward. One recent progress is to construct black holes with “inner extremal horizon”, with zero surface gravity, and therefore avoids the classical blueshift instability. Unfortunately this does not seem to avoid the Aretakis instability (classical perturbation does not decay) [23, 24, 25, 26]. In addition, semi-classically the inner extremal horizon remains unstable [27]. This way of obtaining a remnant is therefore quite challenging conceptually, but clearly more studies are required, as the properties of inner horizon of a black hole are still unclear in general [28], and the possibility that black holes can asymptote towards some extremal configurations and thus become an effective remnant cannot yet be ruled out [29]. Note that extremal black hole is not the only possible end state, as some regular black holes can evolve into a horizonless object [30, 31, 32].
@@
Another possibility is to consider quantum gravitational effect that kicks in when the black hole is sufficiently small. Note the subtle difference: in the contexts of regular black holes, while the classical singularity could be removed by some kind of quantum gravity effect, remnants are the result of the existence of extremal state. The inner horizon — which would not be there without the quantum gravitational effect regularizing the singularity — exists already when the black hole is still astronomically large. Thus, these extremal remnants are not directly caused by any quantum gravitational correction, only indirectly so [^7] Admittedly, what qualifies as “direct” is rather subjective.. A more direct influence from quantum gravity can happen at or near the Planck scale, without necessarily modifying the black hole solution when $M$ is still large, and this can be studied phenomenologically by employing the “generalized uncertainty principle” (GUP).
@@
where $L_{p}$ is the Planck length, and $\alpha$ is a dimensionless parameter, typically taken to be of order unity. We will first discuss the case in which $\alpha>0$. From Eq.(2) we can obtain the following inequality:
@@
Hawking temperature is modified under GUP. For an asymptotically flat Schwarzschild black hole this modification was obtained heuristically in [33] by considering the characteristic energy of the Hawking particle as $E=pc$ and by identifying $\Delta x$ with the horizon scale $\Delta x\sim 2GM/c^{2}$. This gives (upon multiplying with a normalization constant prefactor $1/2\pi$ in order to obtain the correct temperature when $\alpha\to 0$):
@@
Note that the minus sign before the square root is chosen (instead of the positive one; c.f. Eq.(3)) so that the correct Schwarzschild temperature is obtained in the $\alpha\to 0$ limit. The problem with such a heuristic treatment is that we cannot be sure of its validity. After all, while the method works for the Schwarzschild case, it does not lead to the correct known temperature [^9] By “known temperature” we mean the expression obtained using standard methods like Hawking’s original Bogoliubov transformation, the Euclidean Wick rotation method, quantum anomaly calculations, or even tunneling method. They all give the same temperature expression. for the Kerr or Reissner-Nordström black hole [34] (though intriguingly, how it fails may be related to black hole thermodynamics [34]), so how could we be confident that it works when there is a GUP correction? For other criticisms for heuristic treatments of GUP, see [35, 109, 37].
@@
where $M_{p}$ denotes the Planck mass. The fact that there is a nonzero temperature yet the black hole no longer evaporates (otherwise the square root term in Eq.(5) becomes imaginary) could be a sign that modifying the Hawking temperature without modifying the Schwarzschild metric may be problematic. Furthermore, if we look at the temperature of a Reissner-Nordström black hole in standard general relativity, there is also a square root involved, namely a factor of $\sqrt{M^{2}-Q^{2}}$ (in Planck units). In this case the temperature tends to zero as $M\to Q$, so the $M=Q$ state is an extremal black hole. Still, even then, there are considerable interests in the possibility of a naked singularity formation. That is, whether some kind of perturbation can bring $Q$ to exceed $M$. Crucially, note that $\sqrt{M^{2}-Q^{2}}\notin\mathbb{R}$ does not imply there is no state such that $Q>M$, only that such a state would no longer be a black hole. Therefore, in the GUP case discussed above, it is actually not necessary that the $T\neq 0$ remnant “ceases radiating” because the temperature becomes imaginary. Rather, precisely because $T\neq 0$, there is a risk that it may drive the evolution further beyond a black hole stage into another horizonless configuration, if not a naked singularity. Whether this is the case would require a careful investigation.
Of course, the whole idea of a classical, well-defined, metric tensor might already fail near the Planck scale, so it is not clear if any modified metric is trustworthy (worse still, different metrics can be obtained by relying on different heuristic “derivations” [109]). Nevertheless, we note that the entropy of the black hole, which can be obtained by integrating the first law of black hole thermodynamics, can be shown to satisfy (returning to Planck units hereinafter)
@@
Surprisingly, the problem of the final non-evaporating but nonzero temperature black hole can be avoided. This is due to another effect of GUP that can prevent black hole evaporation even when the black hole is still astronomically large (and therefore when the temperature remains largely unmodified), just after a scrambling time [38]. If correct, this renders the aforementioned Planckian size black hole remnant irrelevant except for black holes that were formed to be microscopic in the beginning. For a Schwarzschild black hole, the scrambling time [39, 40] is $r_{S}\log(r_{S}/L_{p})$, where $r_{S}=2M$ denotes the horizon; it is the time scale for information to be scrambled over the horizon after being dropped into a black hole. Note that this time scale is far shorter than $\sim r_{S}^{3}$, the lifetime of a black hole. Thus, if a large black hole stops evaporating after the scrambling time, then it becomes a macroscopic remnant and the peculiarity mentioned above never arises. Intuitively this result can be appreciated as follows. First, we recall the heuristic derivation of Hawking temperature above, in which the Hawking temperature scale is roughly set by $\Delta p$, and by the uncertainty principle $\Delta x\sim 1/\Delta p$, we can trace $\Delta x$ back to the horizon scale, i.e., Hawking radiation comes from the region around the black hole. However, with GUP, it can be shown that beyond the scrambling time, $\Delta p\gg\alpha/L_{p}^{2}$, which in turn implies that $\Delta x\gg r_{S}$, which cannot be interpreted as particle creation by the black hole [38]. See also [41], in which the same conclusion was reached using a string theoretic approach. Note that if correct, this also avoids the firewall problem since the black hole never becomes small enough to get past the Page time scale. The remnant is essentially a classical black hole. To contrast with this scrambling time scale remnant, in another recent approach, it is argued that a black hole stops evaporating after it has emitted about half of its initial mass, due to “memory burden effect” [42], in which a system with high capacity of information storage (such as a black hole) is stabilized by the load of information it carries. Thus we see that it is not necessarily the case that remnants are microscopic.
@@
It is almost poetic that a quantum gravitational correction like the one in GUP, which we usually think can only be important when the black hole is small, can give rise to a drastic deviation from semi-classical picture even when the black hole is still huge (and even renders the black hole effectively classical with no Hawking radiation). Yet, this is not the first time we have seen such an unexpected behavior. For another example, consider the Chandrasekhar limit of a white dwarf star. White dwarfs are stable because the electron degenerate pressure counteracts gravitational collapse. This pressure is inherently quantum mechanical in nature, it arises from the jittering motion (large $\Delta p$) of the particles inside the star, which increases as the star becomes smaller (small $\Delta x$). Once GUP is taken into account, it seems that the Chandrasekhar limit is destroyed [43, 44]. Dynamical stability analysis can restore an upper mass limit for white dwarfs for a certain range of the GUP parameter [45], which explains why in astrophysical observation we do not observe arbitrarily large white dwarfs. Still, here too we see that a supposedly small quantum gravity effect can lead to big changes in physics well outside the quantum gravity regime.
@@
In [46] I considered Hawking evaporation under the assumption that the GUP parameter could be negative. In such a scenario there is no longer a minimum mass, and so the black hole can continue to evaporate indefinitely. However, its lifetime turns out to be infinite, so we still have a remnant that asymptotes to zero mass. There is, however, a peculiar property — the zero mass black hole (or rather, no more black hole!) has a finite nonzero temperature in the negative GUP parameter case. This is as weird, if not weirder than, the positive GUP parameter case in which the non-zero temperature remnant supposedly does not radiate. Fortunately, since this bizarre state cannot be reached in a finite time, it is not physically relevant as the end state of an evaporating black hole [^11] There is still one subtle issue: a zero mass black hole should be easily produced by quantum fluctuations, so why don’t we see them? The reason may have to do with their nonzero temperature, which in a consistent quantum gravity treatment may contribute to a nonzero effective energy-momentum tensor and this then gives it an effective mass after all.. See also [47] for more discussions.
@@
So far we have discussed extremal black holes and the GUP approaches that led to remnants. In recent years, another promising possibility has come to light: a white hole remnant, which is a scenario supported by loop quantum gravity [48, 49, 50, 51, 52, 53, 54], see the recent review [55] for details (and [56] for a nice summary). The basic idea is simple: the static Schwarzschild solution does not distinguish the direction of time, and indeed in the maximally extended spacetime we see that there are black hole and white hole, described by the same metric tensor. Classically, we do not usually care about the white hole part because it is unphysical (in the sense that being a time-reversed version of black hole, a white hole cannot be produced by gravitational collapse). However, quantum mechanically one should consider a superposition of both black and white holes, and the possible tunneling from the black hole state to the white hole one. This probability becomes significant as the black hole becomes sufficiently small. In this scenario, a black hole with initial mass $M$ evaporates with lifetime that scales like $M^{3}$ as in the standard Hawking picture. However, before it completely evaporates away, it tunnels into the white hole state, which then becomes a remnant that slowly radiates away. The time scale for this diffuse emission is about $M^{4}$ [54]. Such a white hole remnant does not live indefinitely. However, like the more traditional black hole remnant it can conceivably still solve the information paradox, as we will discuss in the next section.